\begin{thebibliography}{99}

\bibitem{electron_g} D. Hanneke, S. Fogwell and G. Gabrielse, Phys. Rev. Lett. \textbf{100}, 120801 (2008).

\bibitem{muon_SM} M. Davier, A. Hoecker, B. Malaescu, C. Z. Yuan and Z. Zhang, arXiv:0908.4300 [hep-ph].

\bibitem{muon_g} G. W. Bennett \emph{et al} [Muon g-2 Collaboration], Phys. Rev. D \textbf{73}, 072003 (2006).

%\cite{Frampton:1994rk}
\bibitem{Frampton:1994rk}
  P.~H.~Frampton and T.~W.~Kephart,
  %``Simple nonAbelian finite flavor groups and fermion masses,''
  Int.\ J.\ Mod.\ Phys.\  A {\bf 10} (1995) 4689
  [arXiv:hep-ph/9409330].
  %%CITATION = IMPAE,A10,4689;%%

%\cite{Frampton:2007et}
\bibitem{Frampton:2007et}
  P.~H.~Frampton and T.~W.~Kephart,
  %``Flavor Symmetry for Quarks and Leptons,''
  JHEP {\bf 0709} (2007) 110
  [arXiv:0706.1186 [hep-ph]].
  %%CITATION = JHEPA,0709,110;%%

%\cite{Frampton:2008bz}
\bibitem{Frampton:2008bz}
  P.~H.~Frampton, T.~W.~Kephart and S.~Matsuzaki,
  %``Simplified Renormalizable $T^{'}$ Model for Tribimaximal Mixing and Cabibbo
  %Angle,''
  Phys.\ Rev.\  D {\bf 78} (2008) 073004
  [arXiv:0807.4713 [hep-ph]].
  %%CITATION = PHRVA,D78,073004;%%

\bibitem{MuonAnomalousG-2}
  E.~Ma and M.~Raidal,
  %``Neutrino mass, muon anomalous magnetic moment, and lepton flavor
  %nonconservation,''
  Phys.\ Rev.\ Lett.\  {\bf 87} (2001) 011802
  [Erratum-ibid.\  {\bf 87} (2001) 159901]
  [arXiv:hep-ph/0102255];
  T.~W.~Kephart and H.~Pas,
  %``Muon anomalous magnetic moment in string inspired extended family
  %models,''
  Phys.\ Rev.\  D {\bf 65} (2002) 093014
  [arXiv:hep-ph/0102243];
  Z.~H.~Xiong and J.~M.~Yang,
  %``Muon anomalous magnetic moment in technicolor models,''
  Phys.\ Lett.\  B {\bf 508} (2001) 295
  [arXiv:hep-ph/0102259];
  Z.~z.~Xing,
  %``Nearly bi-maximal neutrino mixing, muon g-2 anomaly and
  %lepton-flavor-violating processes,''
  Phys.\ Rev.\  D {\bf 64} (2001) 017304
  [arXiv:hep-ph/0102304];
  T.~Ibrahim, U.~Chattopadhyay and P.~Nath,
  %``Constraints on explicit CP violation from the Brookhaven muon g-2
  %experiment,''
  Phys.\ Rev.\  D {\bf 64} (2001) 016010
  [arXiv:hep-ph/0102324];  J.~R.~Ellis, D.~V.~Nanopoulos and K.~A.~Olive,
  %``Combining the muon anomalous magnetic moment with other constraints on  the
  %CMSSM,''
  Phys.\ Lett.\  B {\bf 508} (2001) 65
  [arXiv:hep-ph/0102331];  X.~Calmet, H.~Fritzsch and D.~Holtmannspotter,
  %``The anomalous magnetic moment of the muon and radiative lepton decays,''
  Phys.\ Rev.\  D {\bf 64} (2001) 037701
  [arXiv:hep-ph/0103012];  K.~Choi, K.~Hwang, S.~K.~Kang, K.~Y.~Lee and W.~Y.~Song,
  %``Probing the messenger of supersymmetry breaking by the muon anomalous
  %magnetic moment,''
  Phys.\ Rev.\  D {\bf 64} (2001) 055001
  [arXiv:hep-ph/0103048];  S.~Rajpoot,
  %``Muon anomalous magnetic moment in models with singlet fermions,''
  arXiv:hep-ph/0103069;  C.~A.~de S.Pires and P.~S.~Rodrigues da Silva,
  %``Scalar scenarios contributing to (g-2)(mu) with enhanced Yukawa
  %couplings,''
  Phys.\ Rev.\  D {\bf 64} (2001) 117701
  [arXiv:hep-ph/0103083];  E.~O.~Iltan and H.~Sundu,
  %``Anomalous magnetic moment of muon in the general two Higgs doublet
  %model,''
  Acta Phys.\ Slov.\  {\bf 53} (2003) 17
  [arXiv:hep-ph/0103105];  S.~Baek, P.~Ko and H.~S.~Lee,
  %``Muon anomalous magnetic moment, B --> X/s gamma and dark matter  detection
  %in the string models with dilaton domination,''
  Phys.\ Rev.\  D {\bf 65} (2002) 035004
  [arXiv:hep-ph/0103218];  M.~Raidal,
  %``Enhancement of radiatively induced magnetic moment form-factors of  muon:
  %An effective Lagrangian approach,''
  Phys.\ Lett.\  B {\bf 508} (2001) 51
  [arXiv:hep-ph/0103224];  A.~Dedes and H.~E.~Haber,
  %``A light Higgs boson explanation for the g-2 crisis,''
  arXiv:hep-ph/0105014;  E.~O.~Iltan,
  %``The noncommutative effects on the dipole moments of fermions in the
  %standard model,''
  JHEP {\bf 0305} (2003) 065
  [arXiv:hep-ph/0304097];  H.~Chavez, C.~N.~Ferreira and J.~A.~Helayel-Neto,
  %``Physics beyond the standard model: Focusing on the muon anomaly,''
  Phys.\ Rev.\  D {\bf 74} (2006) 033006
  [arXiv:hep-ph/0410373];
  %%CITATION = PHRVA,D74,033006;%%
  A.~Mondragon, M.~Mondragon and E.~Peinado,
  %``S_3-flavour symmetry as realized in lepton flavour violating processes,''
  J.\ Phys.\ A  {\bf 41}, 304035 (2008)
  [arXiv:0712.1799 [hep-ph]].
  %%CITATION = JPAGB,A41,304035;%%


  %%%%%%%%%%%%%%%%


\bibitem{FamilySymmetryRefs}
  M.~Schmaltz,
  %``Neutrino oscillations from discrete nonAbelian family symmetries,''
  Phys.\ Rev.\  D {\bf 52} (1995) 1643
  [arXiv:hep-ph/9411383];
  L.~J.~Hall and H.~Murayama,
  %``A Geometry of the generations,''
  Phys.\ Rev.\ Lett.\  {\bf 75} (1995) 3985
  [arXiv:hep-ph/9508296];
  C.~D.~Carone and R.~F.~Lebed,
  %``A hexagonal theory of flavor,''
  Phys.\ Rev.\  D {\bf 60} (1999) 096002
  [arXiv:hep-ph/9905275];
  P.~H.~Frampton and A.~Rasin,
  %``Nonabelian discrete symmetries, fermion mass textures and large  neutrino
  %mixing,''
  Phys.\ Lett.\  B {\bf 478} (2000) 424
  [arXiv:hep-ph/9910522];
  R.~Dermisek and S.~Raby,
  %``Fermion masses and neutrino oscillations in SO(10) SUSY GUT with  D(3) x
  %U(1) family symmetry,''
  Phys.\ Rev.\  D {\bf 62} (2000) 015007
  [arXiv:hep-ph/9911275];
  G.~Altarelli and F.~Feruglio,
  %``Models of neutrino masses and mixings,''
  New J.\ Phys.\  {\bf 6} (2004) 106
  [arXiv:hep-ph/0405048];
  K.~S.~Babu and J.~Kubo,
  %``Dihedral families of quarks, leptons and Higgses,''
  Phys.\ Rev.\  D {\bf 71} (2005) 056006
  [arXiv:hep-ph/0411226];
  N.~Haba and K.~Yoshioka,
  %``Discrete flavor symmetry, dynamical mass textures, and grand
  %unification,''
  Nucl.\ Phys.\  B {\bf 739} (2006) 254
  [arXiv:hep-ph/0511108];
  Y.~Kajiyama, E.~Itou and J.~Kubo,
  %``Nonabelian discrete family symmetry to soften the SUSY flavor problem  and
  %to suppress proton decay,''
  Nucl.\ Phys.\  B {\bf 743} (2006) 74
  [arXiv:hep-ph/0511268];
  C.~Hagedorn, M.~Lindner and R.~N.~Mohapatra,
  %``S(4) flavor symmetry and fermion masses: Towards a grand unified theory  of
  %flavor,''
  JHEP {\bf 0606} (2006) 042
  [arXiv:hep-ph/0602244];;
  T.~Kobayashi, H.~P.~Nilles, F.~Ploger, S.~Raby and M.~Ratz,
  %``Stringy origin of non-Abelian discrete flavor symmetries,''
  Nucl.\ Phys.\  B {\bf 768} (2007) 135
  [arXiv:hep-ph/0611020];
  M.~C.~Chen and K.~T.~Mahanthappa,
  %``CKM and Tri-bimaximal MNS Matrices in a $SU(5) \times ^{(d)}T$ Model,''
  Phys.\ Lett.\  B {\bf 652} (2007) 34
  [arXiv:0705.0714 [hep-ph]];
  M.~Frigerio and E.~Ma,
  %``Common origin of \theta_{13} and \Delta m^2_{12} in a model of neutrino
  %mass with quaternion symmetry,''
  Phys.\ Rev.\  D {\bf 76} (2007) 096007
  [arXiv:0708.0166 [hep-ph]];
  S.~Sen,
  %``Quark Masses in Supersymmetric SU(3)color x SU(3)W x U(1)X Model with
  %Discrete T' flavor symmetry,''
  Phys.\ Rev.\  D {\bf 76} (2007) 115020
  [arXiv:0710.2734 [hep-ph]];
  G.~Altarelli,
  %``Lectures on Models of Neutrino Masses and Mixings,''
  arXiv:0711.0161 [hep-ph];
  N.~Kifune, J.~Kubo and A.~Lenz,
  %``Flavor Changing Neutral Higgs Bosons in a Supersymmetric Extension based on
  %a $Q_{6}$ Family Symmetry,''
  Phys.\ Rev.\  D {\bf 77} (2008) 076010
  [arXiv:0712.0503 [hep-ph]];
  M.~Honda and M.~Tanimoto,
  %``Deviation from tri-bimaximal neutrino mixing in A4 flavor symmetry,''
  Prog.\ Theor.\ Phys.\  {\bf 119} (2008) 583
  [arXiv:0801.0181 [hep-ph]];
  G.~Altarelli, F.~Feruglio and C.~Hagedorn,
  %``A SUSY SU(5) Grand Unified Model of Tri-Bimaximal Mixing from A4,''
  JHEP {\bf 0803} (2008) 052
  [arXiv:0802.0090 [hep-ph]];
  F.~Plentinger, G.~Seidl and W.~Winter,
  %``Group Space Scan of Flavor Symmetries for Nearly Tribimaximal Lepton
  %Mixing,''
  JHEP {\bf 0804} (2008) 077
  [arXiv:0802.1718 [hep-ph]];
  C.~Luhn,
  %``Discrete Anomalies of Binary Groups,''
  Phys.\ Lett.\  B {\bf 670} (2009) 390
  [arXiv:0807.1749 [hep-ph]];
  P.~H.~Frampton and S.~Matsuzaki,
  %``Prediction for $\Gamma(H\to \tau^+\tau^-)/ \Gamma(H\to \mu^+\mu^-)$ from
  %Non-Abelian Flavor Symmetry,''
  Mod.\ Phys.\ Lett.\  A {\bf 24} (2009) 2081
  [arXiv:0810.1029 [hep-ph]];
  D.~A.~Eby, P.~H.~Frampton and S.~Matsuzaki,
  %``Predictions of Neutrino Mixing Angles in a T'Model,''
  Phys.\ Lett.\  B {\bf 671} (2009) 386
  [arXiv:0810.4899 [hep-ph]];
  F.~Bazzocchi and S.~Morisi,
  %``S4 as a natural flavor symmetry for lepton mixing,''
  Phys.\ Rev.\  D {\bf 80} (2009) 096005
  [arXiv:0811.0345 [hep-ph]];
  P.~H.~Frampton and S.~Matsuzaki,
  %``$T^{'}$ Predictions of PMNS and CKM Angles,''
  Phys.\ Lett.\  B {\bf 679} (2009) 347
  [arXiv:0902.1140 [hep-ph]];
  D.~A.~Eby, P.~H.~Frampton and S.~Matsuzaki,
  %``Analysis of Quark Mixing Using Binary Tetrahedral Flavor Symmetry,''
  Phys.\ Rev.\  D {\bf 80} (2009) 053007
  [arXiv:0907.3425 [hep-ph]];
  M.~C.~Chen, K.~T.~Mahanthappa and F.~Yu,
  %``A Viable Randall-Sundrum Model for Quarks and Leptons with T' Family
  %Symmetry,''
  [arXiv:0907.3963 [hep-ph]];
  %%CITATION = ARXIV:0907.3963;%%
  M.~C.~Chen and K.~T.~Mahanthappa,
  %``Group Theoretical Origin of CP Violation,''
  Phys.\ Lett.\  B {\bf 681}, 444 (2009)
  [arXiv:0904.1721 [hep-ph]];
  %%CITATION = PHLTA,B681,444;%%
%\cite{Feruglio:2008ht}
  F.~Feruglio, C.~Hagedorn, Y.~Lin and L.~Merlo,
  %``Lepton Flavour Violation in Models with A4 Flavour Symmetry,''
  Nucl.\ Phys.\  B {\bf 809}, 218 (2009)
  [arXiv:0807.3160 [hep-ph]];
  %%CITATION = NUPHA,B809,218;%%
%\cite{Feruglio:2007uu}
  F.~Feruglio, C.~Hagedorn, Y.~Lin and L.~Merlo,
  %``Tri-bimaximal neutrino mixing and quark masses from a discrete flavour
  %symmetry,''
  Nucl.\ Phys.\  B {\bf 775}, 120 (2007)
  [arXiv:hep-ph/0702194].
  %%CITATION = NUPHA,B775,120;%%



\bibitem{LEPII} R. Barate \textit{et al.} [LEP Collaborations], Phys. Lett. B \textbf{565}, 61 (2003).


%%%%%%%%%%%%%%%%

%%%%g-2 model refs






\end{thebibliography}

\documentclass[preprint,superscriptaddress,eqsecnum,nofootinbib]{revtex4}
\usepackage{graphicx,amssymb,amsbsy,bm,amsmath,psfrag,dsfont}
\usepackage{slashed}
\usepackage{hyperref} \usepackage[dvips]{color}

\begin{document}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\no}{\noindent}

\newcommand{\la}{\lambda}
\newcommand{\si}{\sigma}
\newcommand{\vk}{\vec{k}}
\newcommand{\vx}{\vec{x}}
\newcommand{\om}{\omega}
\newcommand{\Om}{\Omega}
\newcommand{\ga}{\gamma}
\newcommand{\Ga}{\Gamma}
\newcommand{\gaa}{\Gamma_a}
\newcommand{\al}{\alpha}
\newcommand{\ep}{\epsilon}
\newcommand{\app}{\approx}
\newcommand{\uvk}{\widehat{\bf{k}}}
\newcommand{\OM}{\overline{M}}

\title{Electron and Muon $g-2$ Contributions from the $T'$ Higgs Sector}
\author{Chiu Man Ho} \email{chiuman.ho@vanderbilt.edu}
  \affiliation{Department of
  Physics and Astronomy, Vanderbilt University, Nashville, Tennessee
  37235, USA}
\author{Thomas W. Kephart}\email{tom.kephart@gmail.com}
\affiliation{Department of
  Physics and Astronomy, Vanderbilt University, Nashville, Tennessee
  37235, USA}

\date{\today}

\begin{abstract}
We study the experimental constraints from electron and muon
$g-2$ factors on the Higgs masses and Yukawa couplings in the $T'$ model, and thereby show that the
discrepancy between the standard model prediction and experimental value of muon $g-2$ factor can be
easily accommodated.
\end{abstract}

\maketitle

\section{Introduction}

The electron anomalous magnetic moment has been measured  to an extremely high precision and agrees with the theoretical prediction calculated from
the standard model (SM) \cite{electron_g}, with the result
\bea
\label{e_constraint}
\Delta a_e= |a_e^{\rm{SM}}-a_e^{\rm{Expt}}| < 1\times 10^{-10}.
\eea
On the other hand, the most recent theoretical calculation of the muon anomalous magnetic moment gives \cite{muon_SM}:
\bea
a_\mu^{\rm{SM}} = (11 659 183.4 \pm 4.9) \times 10^{-10},
\eea
where the errors are dominated by the hadronic contribution.
The corresponding most updated experimental value is \cite{muon_g}:
\bea
a_\mu^{\rm{Expt}} = (11 659 208.0 \pm 5.4 \pm 3.3) \times 10^{-10}.
\eea
This implies that $a^{\rm{SM}}_\mu$ differs from $a_\mu^{\rm{Expt}}$ by $3.1 \sigma$, and suggests that a contribution beyond standard model
 may be required. As we will show, this discrepancy between the theoretical and experimental values can be easily accommodated
in the $T'$ model \cite{Frampton:1994rk,Frampton:2007et,Frampton:2008bz} due to the existence of a new and unique Higgs coupling to the muon. While many authors have developed models that resolve this discrepancy \cite{MuonAnomalousG-2},  only a few have  invoked a discrete flavor symmetry.

\section{ Higgs Contributions to $g-2$ Factors in the $T'$ Model}

The $T'$ model \cite{Frampton:1994rk,Frampton:2007et,Frampton:2008bz} relates quarks and electrons through a discrete
flavor symmetry, the binary tetrahedral group $T'$, whose irreducible representations are three singlets, three doublets and a triplet.
The  renormalizable $T'$ model has led to successful predictions of the tribimaximal neutrino mixing matrix as well as the Cabibbo angle \cite{Frampton:2007et,Frampton:2008bz}.
More details about the $T'$ model, its variants and other related models can be found in the literature \cite{FamilySymmetryRefs}.

In the $T'$ model, electrons and muons couple to the different components of the triplet Higgs $H'_3$ through the
interaction terms $Y_e\, \bar{e}\, H'_{3,\,e} \, e$ and $Y_\mu\, \bar{\mu}\, H'_{3,\,\mu}\, \mu$. To compute the contribution of a virtual
Higgs to the electron and muon $g-2$ factors, we need to study its contribution to the electron/muon-photon vertex. For $f=e,\,\mu$, the
vertex function is given by

\bea
&& -i\, e\, \bar{u}(p')\, \Lambda^{\nu}_f (p',p) \,u(p) \nonumber \\
&=& (-i\,e) (-i \,Y_f)^2 \,\int\,\frac{d^4\,k}{(2\pi)^4}\,\bar{u}(p')\,
\frac{i}{k^2-M^2_{H_f}+i\epsilon}\,\frac{i\,(\slashed{p'}-\slashed{k}+m)}{(p'-k)^2-m^2+i\epsilon}\,
\gamma^{\nu}\,\frac{i\,(\slashed{p}-\slashed{k}+m_f)}{(p-k)^2-m_f^2+i\epsilon} \,u(p). \nonumber \\
\eea
where $\bar{u}(p')$ and $u(p)$ are the spinors obeying the equation of motions
$\bar{u}(p')(\slashed{p'}-m_f)=(\slashed{p}-m_f)u(p)=0$, and $M_{H_f}$ is the mass of the Higgs which couples to the electron or muon whose mass
is denoted by $m_f$.

After some calculations, we obtain
\bea
\bar{u}(p')\, \Lambda^{\nu}_f (p,p') \,u(p)= F_f(q^2)\, \bar{u}(p')\, \frac{i\,\sigma^{\nu\al}\,q_\al}{2\,m_f}\,\,u(p)+ \cdots\,,
\eea
where $F_f(q^2)$ is the form factor associated with the electron or muon, and $\sigma^{\nu\al}=\frac{i}{2}[\ga^\nu,\ga^\al]$. The
contributions from the $T'$ Higgs sector to electron or muon anomalous magnetic moment is given by
\bea
\Delta a_f&=&\Delta \left(\,\frac{g_f-2}{2}\,\right)=F_f(q^2=0)\\
&=& \frac{Y_f^2}{8\pi^2}\,\frac{m_f^2}{M^2_{H_f}}\,\int_{0}^{1}\, dx\, \frac{(1-x^2)(1-x)}{x+(1-x)^2\,\frac{m_f^2}{M^2_{H_f}}}\,.
\eea
For $m_f \ll M_{H_f}$, which is likely to be the case, there is a logarithmic divergence in the above integral as $x\rightarrow 0$. This divergence
can be extracted by setting $1-x \rightarrow 1$ and $1-x^2 \rightarrow 1$ in the integrand. As a result, we obtain
\bea
\label{Delta_a}
\Delta a_f \approx \frac{Y_f^2}{4 \pi^2}\,\left(\,\frac{m_f}{M_{H_f}}\,\right)^2\,\ln \left(\,\frac{M_{H_f}}{m_f}\, \right)\,.
\eea
Note that for a given value of $Y_f$,\, $\Delta a_f$ is strictly decreasing when the ratio $M_{H_f}/ m_f$ increases. \\

The condition (\ref{e_constraint}) implies that any combinations of $Y_e$ and $M_{H_e}$ must be such that
\bea
|\Delta a_e| < 1\times 10^{-10},
\eea
which imposes the following constraint
\bea
\label{muon_constraint2}
Y_{e} \,\lesssim\, 21.4 \,\lambda_e \, \frac{M_{H_e}/m_{e}}{\sqrt{\,\ln \left(\,M_{H_e}/m_{e}\,\right)}}\,,
\eea
where $\lambda_e \sim 3 \times 10^{-6} $ is the corresponding electron Yukawa coupling in SM. We required the ratio $M_{H_e}/m_e \gg 1$ when we were deriving (\ref{Delta_a}), but otherwise a free parameter.  To have an assessment on the
allowed range of $Y_e$, we need to have some experimental bounds on $M_{H_e}$. Apparently, we would have hoped that the LEP \cite{LEPII} bound on Higgs mass may help --- due to the non-observation of the  ``Higgs-strahlung" process $e^{+}\,e^{-}\rightarrow H\, Z$ at LEP, a lower bound has been given to the SM Higgs, namely $M_{H_{\textrm SM}} \geq 114.5 \,$ GeV. However, in the $T'$ model, all the Higgs singlets and triplets couple to $Z$. Thus, the LEP bound does not apply directly to any of the masses of the Higgs singlets and triplets. If we simply assume that $M_{H_e} \gtrsim 100 $ GeV, then we  require $Y_e \lesssim 3.5$ in order to satisfy the condition (\ref{e_constraint}) . In this case, the upper bound on the Yukawa coupling $Y_e$ is very loose and any value of $Y_e$ that is perturbatively small would be allowed. \\

For the muon anomalous magnetic moment, the discrepancy between the theoretical and experimental values can be accounted for easily
in the $T'$ model if
\bea
\label{muon_constraint}
\Delta a_\mu \sim |a_\mu^{\rm{SM}}-a_\mu^{\rm{Expt}}| = (24.6 \pm 8.0)\times 10^{-10}\,,
\eea
leading to the constraint
\bea
\label{muon_constraint2}
Y_{\mu} \sim 0.52 \, \lambda_{\mu}\,\frac{M_{H_\mu}/m_{\mu}}{\sqrt{\,\ln \left(\,M_{H_\mu}/m_{\mu}\,\right)}}\,,
\eea
where $\lambda_\mu \sim 0.0006$ is the corresponding muon Yukawa coupling in SM. It is obvious that $Y_\mu \gg \lambda_\mu$, for any choice of
$M_{H_\mu}/m_\mu \gg 1$. For instance, if we assume that $M_{H_\mu} \gtrsim 100 \, $GeV, then in order to satisfy (\ref{muon_constraint2}), we require
$Y_\mu \gtrsim 0.13$.
%In fact, if the mixing between $H'_{3,\,\mu}$ and $H'_{3,\,e}$ is significant, then the LEP bound on Higgs mass will more or less apply to %$M_{H_\mu}$.
%In this case, it is likely that the assumption $M_{H_\mu} \sim 100 \, $GeV is plausible.

\section{Conclusions}

In this article, we have computed the contributions to electron and muon $g-2$ factors from the Higgs sector in the $T'$ model. We then
used the experimental data to constrain the $T'$ model Higgs masses and Yukawa couplings.

If we assume that $M_{H_e} \gtrsim 100 $ GeV, then the upper bound on the electron Yukawa coupling $Y_e$ would be very loose and any value of $Y_e$ consistent with the perturbation theory would be allowed.

Our main result is the demonstration that the discrepancy between the standard model and experimental values of muon anomalous $g-2$ factor can be accounted
for easily in the $T'$ model. Assuming $M_{H_\mu} \gtrsim 100 $ GeV, we found that
the Yukawa coupling $Y_\mu$ should be much larger than the corresponding SM value in order to explain the discrepancy.

\begin{acknowledgments}
We thank Shinya Matsuzaki for useful comments. This work was supported by US DOE
grant DE-FG05-85ER40226.
\end{acknowledgments}

\end{document}